The Stockholders' Equity category of Rausch Company's balance sheet as of December 31, 2017, appeared as follows:

Preferred stock, $100 par, 8%, 2,000 shares
   issued and outstanding $200,000
Common stock, $10 par, 40,000 shares
  issued and outstanding 400,000
Additional paid-in capital 500,000
Total contributed capital $1,100,000
Retained earnings 900,000
    Total stockholders' equity $2,000,000

The notes to the financial statements indicate that dividends were not declared or paid for 2015 or 2016. Rausch wants to declare a dividend of $118,000 for 2017.

Required:

Determine the total and the per-share amounts that should be declared to the preferred and common stockholders under the following assumptions:

  1. The preferred stock is noncumulative, nonparticipating.
  2. The preferred stock is cumulative, nonparticipating.

Round per share amounts to the nearest cent.

Total Preferred Dividends Per Share Amount Total Common Dividends Per Share Amount
1. $ $ $ $
2. $
Answer #1

Dividend distribution

Total preferred dividend Per share amount Total Common dividend Per share amount
1. 200000*8% = 16000 16000/2000 = 8 118000-16000 = 102000 102000/40000 = 2.55
2. 16000*3 = 48000 48000/2000 = 24 118000-48000 = 70000 70000/40000 = 1.75